OpenStudy (anonymous):

if a student goes to school at 2.5 km/hour, he reaches 6 mintues late. If he travels at 3 km/hour, he reaches the school 10 miutes early. what is the distance from his house to school

OpenStudy (anonymous):

lets convert kmph to meters per minute. 2500/60 = 41.66 meters per minute 3000/60 = 50 meters per minute

OpenStudy (anonymous):

41.66(t+6) = 50(t-10)

OpenStudy (anonymous):

t = 90 minutes

OpenStudy (anonymous):

there for distance is 50(90-10) = 4000 meters = 4km